OpenStudy (anonymous):

A truck moves 60 kilometers east from point A to point B. At point B, it turns back west and stops 15 kilometers away from point A. What are the total distance and total displacement of the truck?

OpenStudy (jt950):

Total distance = 45+60 km Total displacement =60-45 km